# Forest Proximities for Time Series

### v2

@@ Abstract
RF-GAP has recently been introduced as an improved random forest proximity measure. In this paper, we present PF-GAP, an extension of RF-GAP proximities to proximity forests, an accurate and efficient time series classification model. We use the forest proximities in connection with Multi-Dimensional Scaling to obtain vector embeddings of univariate time series, comparing the embeddings to those obtained using various time series distance measures. We also use the forest proximities alongside Local Outlier Factors to investigate the connection between misclassified points and outliers, comparing with nearest neighbor classifiers which use time series distance measures. We show that the forest proximities may exhibit a stronger connection between misclassified points and outliers than nearest neighbor classifiers.
Index Terms:
Proximity Forests, time series classification, forest proximities, time series outlier detection, time series embedding methods.
I Introduction
@@
Despite the utility of random forests, including random forest proximities, random forests do not always perform well on time series data [7]. Proximity forests were recently introduced to fill this gap by extending an RF-like algorithm to time-series data. However, proximities for this new forest algorithm have not been previously explored. We introduce forest proximities for proximity forests to extend the applicability of proximity forests for time series data in much the same way as forest proximities extend the applicability of random forests for other data types. Due to their geometry-preserving capabilities, we specifically extend the RF-GAP proximities to proximity forests, and call the resulting proximities PF-GAP.
While there are many potential applications of PF-GAP, we focus our demonstrations on two, namely obtaining vector embeddings for time series visualization and time series outlier detection. We show that using PF-GAP outperforms other approaches in these tasks that use different, common time series distance measures.
II Background
II-A Proximity Forests
Proximity Forests were proposed in 2018 as a comparatively efficient and accurate time series classification model [7]. Early experiments by its creators have shown that proximity forests generally outperform classifiers such as BOSS-VS [8], WEASEL [9], and ElasticEnsemble [10] in terms of accuracy, and that the accuracy is generally comparable to classifiers such as COTE [11].
@@ II-A Proximity Forests
Initially, the exemplars are randomly sampled, and the collection of sampled exemplars is called a candidate split. A total of $r$ candidate splits are randomly generated, where $r$ is a user-defined hyperparameter. The quality of each candidate split is assessed by measuring the Gini purity of the potential child nodes. As the $r$ parameter is increased, the model has a better chance of obtaining the ideal splits at each node. However, increasing the $r$ parameter also increases the computational resources required to train the proximity tree.
@@
II-B Random Forest Proximities
@@
${}p_{GAP}(i,j)=\dfrac{1}{|S_{i}|}\sum_{t\in S_{i}}\dfrac{c_{j}(t)\cdot I(j\in J_{i}(t))}{|M_{i}(t)|}.$ (2)
@@
II-C Time Series Outlier Detection
@@
III Methods
III-A PF-GAP and Pairwise Dissimilarity
@@
${}P(i,j)=\dfrac{1}{2}\left(p_{GAP}(i,j)+p_{GAP}(j,i)\right).$ (3)
@@
We use multidimensional scaling (MDS) to obtain vector embeddings for time series by first computing the pairwise dissimilarity $d_{ij}$, then applying MDS (either metric or non-metric). We note that RF proximities have also been embedded using other methods [18, 6, 19].
Other methods exist for computing pairwise dissimilarities for time series. We consider any symmetric, positive-definite distance measure to be suitable for obtaining dissimilarity, with the distance between observations $i$ and $j$ defining $d_{ij}$. Therefore, we will compare embeddings obtained using $p_{GAP}$ to embeddings obtained using each of the nine time series distance measures named in Section II-A. We should note, however, that while the embeddings obtained using $p_{GAP}$ and a particular time series distance measure are comparable in the sense that the embeddings make use of pairwise dissimilarity, the method using the time series distance measure is unsupervised, whereas the method using $p_{GAP}$ is supervised.
III-B Outlier Detection
@@
${}\sum_{j\in\text{class(i)}}\dfrac{N}{P(i,j)^{2}},$ (6)
@@
In our experiments on outlier detection, we will use LOF for easier comparisons with commonly used dissimilarities for time series. For a given dataset, we compute pairwise distances for each of the nine time series distance measures mentioned previously. For each distance measure, we create a $1$-nearest neighbor classifier, using the classifier to obtain predictions for each time series. Next, we use the pairwise distances with LOF (using $5$ nearest neighbors) to predict outliers. Ideally, a time series is considered an outlier if and only if it is misclassified.
IV Experiments
@@
(b) MDS embedding of the GunPoint dataset using pairwise dissimilarities defined by PF-GAP.
(c) MDS embedding obtained in (b), where the point with the highest within-class outlier score is indicated in red.
@@
The materials needed to reproduce our experiments are available on our project website.  [^1] https://sites.google.com/view/forest-proximities We examine three training datasets from the UCR archive [20], namely GunPoint, ItalyPowerDemand, and ArrowHead. While there are many datasets from the UCR archive, these particular datasets are chosen for two reasons. First, due to the relative ease in classification: our second experiment seeks to make a bijective correspondence between outliers and misclassified points, which is well-posed if misclassified points are in the minority. The second reason these particular datasets are chosen is due to their relatively small size, as obtaining all pairwise dissimilarities for all dissimilarities that we compare can become a resource-consuming task for datasets with a higher number or greater length of time series.
Our first experiment seeks to quantify the inter-class separation of time series vector embeddings. As described in Section III, we compute dissimilarities using $p_{GAP}$ and all nine time series distance measures referred to earlier for each dataset. We call the dissimilarities induced by $p_{GAP}$ DGAP. Then, we obtain embeddings using metric and non-metric MDS.
Apart from visual inspection, we quantify the strength of the inter-class separation of our vector embeddings in our experiments using $k$-means clustering. After obtaining the vector embeddings, for a particular dataset, using time series distance measures and PF-GAP, we apply $k$-means with $k$ chosen to be the (known) number of distinct classes within the dataset. An accuracy score is obtained for each embedding, where the predicted labels are assigned based on which cluster a point is assigned by $k$-means, which score we call the $k$-means clustering score. These scores are shown in Table I. The scores are the higher of the $k$-means clustering scores obtained using metric (indicated with (m)) and non-metric MDS. The best values in each column are bold and the runner-up is underlined.
TABLE I: $k$-means clustering scores.
GunPoint ItalyPowerDemand ArrowHead
DGAP 1.0(m) 0.94(m) 0.75(m)
DTW 0.50 0.60(m) 0.58(m)
DDTW 0.52(m) 0.49(m) 0.5(m)
TWE 0.64 0.49(m) 0.67(m)
WDTW 0.48 0.60(m) 0.56(m)
WDDTW 0.52 0.49(m) 0.5(m)
Euclidean 0.48 0.55(m) 0.67(m)
LCSS 0.74(m) 0.57(m) 0.50
MSM 0.64 0.51(m) 0.67(m)
ERP 0.48 0.58(m) 0.25
The $k$-means clustering scores are highest for DGAP for all three datasets. We offer a visual comparison in two cases: the GunPoint dataset (Figure 1) and the ArrowHead dataset (Figure 2). In both figures, the DTW-induced embeddings are compared with those generated by PF-GAP. The results show that DGAP produces greater interclass separation. The figures also depict the embedded point that corresponds to the time series with the highest within-class outlier score according to Equation 6. In both cases, the indicated point is a visual within-class outlier.
(a) MDS embedding of the ArrowHead dataset using pairwise dissimilarity defined by dynamic time warping distances.
(b) MDS embedding of the ArrowHead dataset using pairwise dissimilarities defined by PF-GAP.
(c) MDS embedding obtained in (b), where the point with the highest within-class outlier score is indicated in red.
Figure 2: MDS embeddings of the ArrowHead dataset.
Next, we turn to the task of identifying misclassified points as outliers, as described in Section III. As with the $k$-means scores, we obtain scores for each of the three datasets with each of the 10 dissimilarity measures (including DGAP). If a time series is classified correctly and is labeled an inlier, we consider this to be a true positive. If a time series is misclassified and labeled an outlier, we consider this to be a true negative. We score each classifier/LOF pair using the F1 score. The results are shown in Table II. As with Table I, the best values in each column are bolded and the runner-up is underlined.
TABLE II: F1 scores for each dataset and metric.
GunPoint ItalyPowerDemand ArrowHead
DGAP 1.0 0.99 0.99
DTW 0.76 0.94 0.73
DDTW 0.85 0.80 0.91
TWE 0.93 0.94 0.88
WDTW 0.80 0.94 0.75
WDDTW 0.85 0.80 0.91
Euclidean 0.87 0.92 0.91
LCSS 0.68 0.67 0.62
MSM 0.89 0.92 0.89
ERP 0.89 0.94 0.89
Most of the F1 scores in Table II are reasonably high, although the F1 scores for DGAP are the highest. The high F1 scores for DGAP demonstrate a possible strong tie between points which are misclassified by the proximity forest algorithm and points that may be considered outliers based on the forest proximities.
V Discussion of Limitations
In both experiments, the dissimilarity defined by $p_{GAP}$ outperformed the dissimilarities defined by commonly used time series distance measures in the datasets we selected. Much of the observed success is likely because the $p_{GAP}$ dissimilarities take class labels into account, whereas the dissimilarities that we compare do not.

### v3

@@ Abstract
RF-GAP has recently been introduced as an improved random forest proximity measure. In this paper, we present PF-GAP, an extension of RF-GAP proximities to proximity forests, an accurate and efficient time series classification model. We use the forest proximities in connection with Multi-Dimensional Scaling to obtain vector embeddings of univariate time series, comparing the embeddings to those obtained using various time series distance measures. We also use the forest proximities alongside Local Outlier Factors to investigate the connection between misclassified points and outliers, comparing with nearest neighbor classifiers which use time series distance measures. We show that the forest proximities seem to exhibit a stronger connection between misclassified points and outliers than nearest neighbor classifiers.
keywords:
Proximity Forests, time series classification, random forest proximities, time series outlier detection, time series embedding methods.
1 Introduction
@@
Despite the utility of random forests, including random forest proximities, random forests do not always perform well on time series data [7]. Proximity forests [7] were recently introduced to fill this gap by extending an RF-like algorithm to time-series data. However, proximities for this new forest algorithm have not been previously explored. We introduce a definition of proximities for proximity forests to extend the applicability of proximity forests for time series data in much the same way as forest proximities extend the applicability of random forests for other data types. Due to their geometry-preserving capabilities, we specifically extend the RF-GAP proximities to proximity forests, and call the resulting proximities PF-GAP.
While there are many potential applications of PF-GAP, we focus mainly on time series outlier detection and the connection between outliers and misclassified points. We show that PF-GAP outperforms other approaches that use different, common time series distance measures. We also demonstrate how PF-GAP can be used to obtain vector embeddings in a supervised manner. Our contributions can be summarized as follows: (1) we implement GAP proximities for the proximity forest model; [^1] See https://sites.google.com/view/forest-proximities for code used to produce the experiments. See also https://github.com/KevinMoonLab/PF-GAP/tree/main for PF-GAP source code. (2) we demonstrate the utility of time series proximities in supervised visualization and within-class outlier detection; (3) we conduct an experiment using 64 datasets from the UCR 2018 archive [8] which suggests a stronger connection between outlier time series and misclassified points than with nearest neighbor classifiers.
This paper is organized as follows. Section 2 gives background information on the proximity forest model and random forest proximities. Additionally, the latter part of section 2 describes the meaning of time series outlier detection as it it used in this paper. Section 3 describes the application of GAP proximities to the proximity forest model, distances induced by the proximities, as well as the method by which within-class outlier scores are obtained. Section 4 provides both the main experimental result of the paper as well as a visual result depicting the use of GAP proximities for time series visualization. Section 5 provides a brief discussion of limitations, and we provide concluding remarks in section 6.
2 Background
2.1 Proximity Forests
Proximity Forests were proposed in 2018 as a comparatively efficient and accurate time series classification model [7]. Early experiments by its creators have shown that proximity forests generally outperform time series classifiers such as BOSS-VS [9], WEASEL [10], and ElasticEnsemble [11] in terms of accuracy, and that the accuracy is generally comparable to classifiers such as COTE [12].
@@ 2.1 Proximity Forests
Initially, the exemplars are randomly sampled, and the collection of sampled exemplars is called a candidate split. A total of $r$ candidate splits are randomly generated, where $r$ is a user-specified hyperparameter. The quality of each candidate split is assessed by measuring the Gini purity of the potential child nodes. As the $r$ parameter is increased, the model has a better chance of obtaining the ideal splits at each node. However, increasing the $r$ parameter also increases the computational resources required to train the proximity tree.
@@
2.2 Random Forest Proximities
@@
${}p_{GAP}(i,j)=\frac{1}{|S_{i}|}\sum_{t\in S_{i}}\frac{c_{j}(t)\cdot I(j\in J_{i}(t))}{|M_{i}(t)|}.$ (2)
@@
2.3 Time Series Outlier Detection
@@
3 Methods
3.1 PF-GAP and Pairwise Dissimilarity
@@
${}P(i,j)=\frac{1}{2}\left(p_{GAP}(i,j)+p_{GAP}(j,i)\right).$ (3)
@@
We call the distance measure defined by equation 5 the DGAP distance.
3.2 Outlier Detection
@@
${}\sum_{j\in\text{class(i)}}\frac{N}{P(i,j)^{2}},$ (6)
@@
In our experiments on outlier detection, we use LOF for easier comparisons with commonly used dissimilarities for time series. For a given dataset, we compute pairwise distances for each of the nine time series distance measures mentioned previously. For each distance measure, we create a $1$-nearest neighbor classifier, using the classifier to obtain predictions for each time series. Next, we use the pairwise distances with LOF (using $5$ nearest neighbors) to predict outliers. Ideally, a time series is considered an outlier if and only if it is misclassified.
4 Experiments
We wish to quantify the relationship between outlier time series and misclassified time series. For each selected training dataset in the UCR 2018 archive [8], we train 10 distinct classifiers corresponding to both the DGAP distance as well as the nine time series distances mentioned previously. In tandem, each distance measure is used to compute pairwise distances which are subsequently used with LOF (using five nearest neighbors) to predict outliers. For each of the 10 distances, F1 scores are obtained: we treat a correctly classified point as a true inlier, using the output of the associated LOF model as an inlier/outlier prediction. Thus, a true positive is a correctly classified point which is an inlier, a true negative is a misclassified point labeled an outlier, a false negative is a correctly classified point which is labeled an outlier, and a false positive is a misclassified point which is labeled an inlier. For the DGAP distance, the associated classifier is a PF model with $r=5$ and 11 trees. For the remaining distances, the associated classifiers are 1-NN classifiers which use the respective time series distances. The F1 scores for each dataset and distance measure are shown in Table 1.
A total of 64 datasets from the UCR archive were used, though there are 128 datasets in the archive. Some were removed due to missing data. However, the majority of removals occurred due to the length of time required to train $1$-NN distance classifiers for 9 distinct distance measures: for some datasets, this task appeared to require multiple days of computing using CPU. [^2] This highlights another advantage of the more efficient proximity forest model. Our selection consisted of datasets for which the $1$-NN classifiers could be computed within a total of 18 minutes.
Table 1: F1 scores for each dataset and distance measure
UCR Dataset DTW DDTW TWE WDTW WDDTW Euclidean LCSS MSM ERP DGAP
BirdChicken 0.76 0.8 0.89 0.71 0.82 0.82 0.75 0.82 0.82 1.0
Lightning7 0.77 0.67 0.84 0.85 0.72 0.78 0.78 0.84 0.83 1.0
InsectWingbeatSound 0.47 0.31 0.65 0.58 0.56 0.65 0.67 0.66 0.58 0.99
Adiac 0.74 0.71 0.76 0.74 0.74 0.74 0.08 0.75 0.75 1.0
FreezerSmallTrain 0.88 0.9 0.85 0.88 0.94 0.96 0.86 0.89 0.89 1.0
MedicalImages 0.76 0.68 0.8 0.77 0.68 0.76 0.71 0.77 0.79 1.0
Rock 0.75 0.69 0.79 0.75 0.92 0.75 0.79 0.79 0.79 1.0
CinCECGTorso 0.64 0.73 0.93 0.93 0.99 0.92 0.96 0.92 0.8 1.0
WordSynonyms 0.75 0.77 0.88 0.74 0.81 0.75 0.79 0.86 0.84 1.0
GunPoint 0.76 0.85 0.95 0.8 0.89 0.89 0.68 0.89 0.89 1.0
FreezerRegularTrain 0.86 0.86 0.92 0.87 0.85 0.85 0.75 0.91 0.91 1.0
FaceFour 0.79 0.85 0.93 0.82 0.75 0.77 0.93 0.91 0.88 1.0
Car 0.72 0.8 0.9 0.7 0.86 0.81 0.49 0.89 0.85 0.99
GunPointMaleVersusFemale 0.93 0.86 1.0 0.92 0.87 0.97 1.0 1.0 0.96 1.0
OliveOil 0.9 0.78 0.91 0.92 0.78 0.91 0.29 0.91 0.91 0.98
MiddlePhalanxTW 0.71 0.75 0.71 0.72 0.76 0.72 0.58 0.7 0.69 0.98
SonyAIBORobotSurface2 0.88 0.84 0.92 0.86 0.87 0.85 0.82 0.9 0.92 1.0
FacesUCR 0.89 0.87 0.98 0.89 0.84 0.86 0.85 0.97 0.96 1.0
PowerCons 0.81 0.75 0.95 0.83 0.77 0.96 0.94 0.95 0.93 1.0
ProximalPhalanxTW 0.85 0.85 0.85 0.85 0.85 0.85 0.08 0.85 0.85 0.99
SyntheticControl 0.94 0.76 0.99 0.94 0.74 0.96 0.86 0.98 1.0 1.0
MiddlePhalanxOutlineCorrect 0.83 0.84 0.87 0.83 0.84 0.87 0.79 0.87 0.87 0.99
MiddlePhalanxOutlineAgeGroup 0.85 0.85 0.84 0.85 0.86 0.86 0.43 0.84 0.84 0.99
Chinatown 0.95 0.95 0.95 0.92 0.95 1.0 0.89 0.92 1.0 1.0
CBF 0.97 0.67 1.0 0.97 0.7 0.91 0.95 0.97 1.0 1.0
InsectEPGRegularTrain 0.87 0.75 0.98 0.91 0.65 0.99 0.91 0.98 0.98 1.0
TwoLeadECG 0.81 0.98 0.93 0.84 0.98 0.88 0.73 0.88 0.95 1.0
ItalyPowerDemand 0.94 0.8 0.94 0.94 0.83 0.92 0.67 0.92 0.94 1.0
InsectEPGSmallTrain 0.87 0.85 0.87 0.87 0.64 0.87 0.9 0.87 0.87 1.0
ToeSegmentation2 0.85 0.8 0.97 0.83 0.87 0.86 0.92 0.99 0.93 1.0
BeetleFly 0.89 0.86 0.92 0.89 0.95 0.71 0.86 0.86 0.86 1.0
Herring 0.66 0.65 0.68 0.61 0.7 0.58 0.68 0.65 0.65 0.99
ECGFiveDays 0.79 0.78 0.86 0.79 0.78 0.9 0.85 0.89 0.87 0.93
DistalPhalanxOutlineCorrect 0.85 0.84 0.86 0.85 0.83 0.85 0.51 0.86 0.87 1.0
Fish 0.85 0.9 0.93 0.82 0.91 0.86 0.26 0.92 0.91 1.0
ToeSegmentation1 0.67 0.53 0.73 0.64 0.47 0.62 0.86 0.84 0.75 0.99
Meat 0.97 0.89 0.97 0.96 0.89 0.99 0.5 0.97 0.98 1.0
Trace 0.91 0.94 0.94 0.89 0.95 0.88 0.84 0.91 0.93 1.0
Symbols 0.7 0.83 0.84 0.7 0.68 0.76 0.68 0.83 0.83 1.0
Ham 0.79 0.84 0.89 0.8 0.82 0.9 0.68 0.9 0.91 1.0
GunPointOldVersusYoung 0.91 0.86 1.0 0.89 0.86 0.95 1.0 1.0 0.96 1.0
ShapeletSim 0.67 0.62 0.95 0.79 0.62 0.52 1.0 0.92 0.95 1.0
HouseTwenty 0.79 0.81 0.87 0.92 0.87 0.77 0.96 0.87 0.96 1.0
OSULeaf 0.81 0.91 0.89 0.83 0.92 0.77 0.81 0.88 0.82 1.0
DistalPhalanxTW 0.83 0.83 0.84 0.83 0.83 0.86 0.75 0.86 0.84 1.0
Wine 0.91 0.92 0.95 0.91 0.92 0.94 0.69 0.95 0.95 1.0
UMD 0.81 0.84 0.82 0.79 0.71 0.82 0.65 0.8 0.85 1.0
Fungi 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.0
ProximalPhalanxOutlineCorrect 0.88 0.9 0.88 0.88 0.9 0.89 0.81 0.88 0.89 0.99
Mallat 0.97 0.86 0.96 0.99 0.94 0.99 0.68 0.95 0.97 1.0
SwedishLeaf 0.82 0.9 0.94 0.82 0.89 0.82 0.36 0.9 0.89 1.0
GunPointAgeSpan 0.89 0.83 0.98 0.88 0.84 0.97 0.96 0.98 0.97 1.0
ECG5000 0.9 0.91 0.96 0.91 0.91 0.96 0.73 0.96 0.96 1.0
Plane 0.96 0.97 0.96 0.95 0.96 0.94 0.84 0.95 0.96 1.0
MoteStrain 0.84 0.74 0.86 0.84 0.74 0.85 0.92 0.86 0.86 1.0
Beef 0.39 0.51 0.49 0.41 0.5 0.51 0.34 0.49 0.5 0.97
SonyAIBORobotSurface1 0.88 0.93 0.94 0.88 0.97 0.91 0.71 0.91 0.89 1.0
DistalPhalanxOutlineAgeGroup 0.86 0.85 0.89 0.85 0.85 0.87 0.65 0.88 0.89 0.99
ArrowHead 0.73 0.91 0.89 0.75 0.91 0.91 0.62 0.89 0.89 0.99
DiatomSizeReduction 0.64 0.93 1.0 0.64 0.9 0.97 0.48 1.0 1.0 1.0
Coffee 0.94 0.98 1.0 0.94 0.98 1.0 0.67 1.0 1.0 1.0
ProximalPhalanxOutlineAgeGroup 0.86 0.88 0.87 0.86 0.87 0.88 0.52 0.86 0.86 0.99
BME 0.82 0.89 0.87 0.87 0.85 0.93 0.68 0.89 0.82 1.0
SmoothSubspace 0.91 0.88 0.99 0.92 0.87 0.95 0.5 0.99 1.0 1.0
In all but seven datasets, the DGAP distances obtained the highest F1 scores, so that DGAP strictly outperformed the other distances approximately $89\%$ of the time. In the remaining seven datasets, the DGAP F1 scores were tied with others which obtained scores of $1.0$: therefore, DGAP was never outperformed. For reference, the names of the UCR datasets for which ties occurred for the highest F1 score are as follows:
  • GunPointMaleVersusFemale
  • Chinatown
  • CBF
  • GunPointOldVersusYoung
  • ShapeletSim
  • DiatomSizeReduction
  • Coffee
The high F1 scores for DGAP demonstrate a possible strong tie between points which are misclassified by the proximity forest algorithm and points that may be considered outliers based on the forest proximities. The connection between misclassified points and outliers appears to be stronger than for $1$-NN classifiers using several commonly-used time series distance measures.
This connection can also be visualized, since the distances defined in equation 5 can be coupled with various manifold learning methods to produce a vector space embedding of the time series data. Additionally, since the distances have been computed in a supervised manner, the resulting embedding will tend to provide better class separation when compared with embeddings based on unsupervised distances. In Figure 1, 2-dimensional embeddings obtained using MDS are shown, offering a visual comparison of class separation obtained using distances defined by $p_{GAP}$ as opposed to distances defined in an unsupervised manner. We note, however, that RF proximities have also been embedded using other methods [19, 6, 20, 21].
@@ 4 Experiments
(b) MDS embedding of the GunPoint dataset using pairwise dissimilarities defined by PF-GAP. The point with the highest within-class outlier score is indicated in red and belongs to Class 2.
@@
As previously shown with tabular data, distances defined by $p_{GAP}$ visually reflect the classification strength of the underlying model, and within-class outliers tend to also be visual outliers [2]. Figure 1 appears to suggest that the same desirable properties tend to apply for PF-GAP.
5 Discussion of Limitations
The dissimilarity defined by $p_{GAP}$ outperformed or tied in with the dissimilarities defined by commonly used time series distance measures for the 64 UCR datasets used. The proximity forest classifier has been shown to generally outperform $1$-NN classifiers [7], which is a conceivable contribution to the higher F1 scores for classification/inlier association. However, the F1 scores allow for weak classifiers, provided that the associated distance measure, coupled with LOF, tends to label misclassified points as outliers.